A bar magnet of length 3 cm has points A and B along its distances of 24 cm and 48 cm on the opposite sides .ratio of magnetic fields at these points will be

Grade:12th pass

1 Answers

Arun
6 years ago
B is inversely proportional to one upon (distance)3 this you see in on axial and normal bisector (magnetic field due to bar magnet)Now, Ba/Bb=(48/24)3 =8/1 answer